HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

EXPLORING THE RESOLUTION LIMIT OF PFI-ZEKE PHOTOELECTRON SPECTROSCOPY

\underline{H. HERBURGER} , V. WIRTH U. HOLLENSTEIN, F. MERKT, Department of Chemistry and Applied Biosciences, ETH Zurich, 8093 Zurich, Switzerland

The spectral resolution in pulsed-field-ionisation zero-kinetic-energy (PFI-ZEKE) photoelectron spectroscopy is related to the state selectivity in the diabatic ionisation of the Rydberg states by time-dependent electric fields. The selectivity is determined by the applied electric-field pulse sequence. Hollenstein et al.^{1} used discrete electricfield pulses with increasing field strength in combination with a preceding field pulse of opposite polarity and achieved a spectral resolution of 0.06 cm^{−1}. To improve the resolution further, Harper et al.^{2} recently replaced the sequence of field steps by a linearly increasing field, as used earlier by Reiser et al.^{3}, in combination with a prepulse of opposite polarity and obtained promising results on the PFI-ZEKE photoelectron spectrum of NO and CO_{2}. They called this new method pulsed-ramped field-ionisation (PRFI)-ZEKE photoelectron spectroscopy.

Using a home-built narrow-bandwidth long-pulse laser system (pulse lengths up to 50 ns, bandwidth< 50 MHz) in combination with a pulsed field ramp,^{2,3} we explore the resolution limit of this approach. To avoid overlap of spectral lines, we chose the atoms Ar, Kr, and Xe as test systems and recorded PRFI-ZEKE photoelectron spectra of transitions from the metastable state (np)^{5}((n + 1)s)[3/2]_{2} (^{3}P_{2}) of the neutral rare-gas atom to the (np)^{5} ^{2}P_{3/2} state of the corresponding ion, with n = 3, 4, 5 for Ar, Kr, and Xe, respectively. We present PRFI-ZEKE photoelectron spectra that exhibit both adiabatic and diabatic ionisation dynamics. The resolution of these spectra is improved to 0.05 cm^{−1} which constitutes the resolution record in photoelectron spectroscopy.^{1} The determined ionisation energies are compared with precisely known literature values^{4}. The current resolution is limited by adiabatic field ionisation below n = 190 and by electric stray fields, which affect the field-ionisation dynamics above n = 190.

1U. Hollenstein, R. Seiler, H. Schmutz, M. Andrist, and F. Merkt, J. Chem. Phys. 115, 5461–5469 (2001).

2Oliver J. Harper, Ning L. Chen, Séverine Boyé-Péronne, and Bérenger Gans, Phys. Chem. Chem.

Phys. 24, 2777–2784 (2022).

3G. Reiser, W. Habenicht, K. Müller-Dethlefs, and E. W. Schlag, Chem. Phys. Lett. 152, 119–123 (1988).

4V. L. Sukhorukov, I. D. Petrov, M. Schäfer, F. Merkt, M.-W. Ruf, and H. Hotop, J. Phys. B: At. Mol.

Opt. Phys. 45, 092001 (2012) and references therein.

p-number: p137  Submitted on Wed, 14 Jun 2023 17:39:46 +0200